OpenStudy (anonymous):

find a normal vector at each point(x,y,z) for the cone z^2=x^2+Y^2 for 0<=z<=1

OpenStudy (rational):

normal vector = \(\langle f_x, f_y, -1 \rangle\)

OpenStudy (rational):

find the partials and plugin

OpenStudy (anonymous):

so my answer would be <2x,2y,-1>?

OpenStudy (rational):

\(z^2 = x^2 + y^2 \implies f_x = \dfrac{x}{z}, f_y = \dfrac{y}{z}\) then the normal vector wud be : \(\left< \dfrac{x}{z}, \dfrac{y}{z}, -1\right>\)

OpenStudy (anonymous):

i derrived and im supposed to integrate right? or how did you get x/z

OpenStudy (rational):

\(z^2 = x^2 + y^2\) \(f_x = \dfrac{\partial z}{\partial x} = ?\) \(f_y = \dfrac{\partial z}{\partial y} = ?\)

OpenStudy (anonymous):

what does that mean sorry

OpenStudy (anonymous):

i have been up all night

OpenStudy (rational):

you familiar with taking partial derivatives, right ?

OpenStudy (anonymous):

yes very im just going blank right now

OpenStudy (anonymous):

fx means derrivae f in terms of x

OpenStudy (rational):

yep !

OpenStudy (rational):

\(z^2 = x^2 + y^2\) partially differentiate both sides in terms of "x"

OpenStudy (rational):

what do u get ?

OpenStudy (rational):

\(z^2 = x^2 + y^2\) partially differentiate both sides in terms of "x" : \(2z \dfrac{\partial z}{\partial x} = 2x + 0\)

OpenStudy (rational):

right ?

OpenStudy (anonymous):

why do you have 2z i get the 2x

OpenStudy (anonymous):

is it the (df/dx)(dx/dz) thing?

OpenStudy (rational):

no lol, we're just differentiating implicitly

OpenStudy (rational):

lets try it directly otherwise

OpenStudy (rational):

\(z^2 = x^2 + y^2\) \(z = \sqrt{x^2 + y^2}\)

OpenStudy (rational):

Now find your partials \(f_x\) and \(f_y\)

OpenStudy (anonymous):

wtf i have no idea whats going on

OpenStudy (rational):

your function is : \(z = f(x,y) = \sqrt{x^2+y^2}\) \(f_x = ?\) \(f_y = ?\)

OpenStudy (anonymous):

so f(x,Y)=x+y then

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

for fx i got\[x \div \sqrt{x ^{2+y ^{2}}}\]

OpenStudy (anonymous):

since z=sqrt(x^2+y^2)

OpenStudy (anonymous):

you get x/z

OpenStudy (rational):

do u mean : \(f_x = \dfrac{x}{\sqrt{x^2+y^2}} = \dfrac{x}{z}\) ?

OpenStudy (anonymous):

yeah

OpenStudy (rational):

Yep ! you got it !!

OpenStudy (anonymous):

lol i had a big brain fart

OpenStudy (anonymous):

can you help me with something more complicated?

OpenStudy (rational):

il try, ask..

OpenStudy (anonymous):

calculate the surface area of the part of the paraboloid z=6-x^2-y^2 above the plane with equation z=2

OpenStudy (rational):

straightforward problem

OpenStudy (rational):

surface area : \[\iint \limits_R \sqrt{f_x^2 + f_y^2 + 1} ~dA\]

OpenStudy (anonymous):

z=6-r^2

OpenStudy (anonymous):

oh i know that lol but my limits of integration

OpenStudy (rational):

for gettign the limits, sketch the region quick

OpenStudy (anonymous):